International Products, Inc. has ordered 20,000 leather coats from Argentina for delivery in 12 months. The contracted cost of a coat is 139 pesos. International Products will pay for the coats upon delivery. The current indirect exchange rate is $1 for 1.2508 pesos. The anticipated inflation rate is 4.2% in the United States and 7.8% in Argentina. In U.S. dollars, how much will the

20,000 leather coats cost International Products at delivery?
Answer #1

Forward rate Spot Rate*(1+Inflation Rate Argentina)/(1+Inflation rate dollar)

= 1.2508(1+7.8%)/(1+4.2%)

= 1.2940 pesos/Dollar

Amount to be paid = 20,000*139 = Pesos 2,780,000

Hence, cost to International Products in Dollars = 2,780,000/1.2940

= $2,148,377.13